abc company has 1/4 computer than employees, if they have 80 employee and want to employ 5% more, how many computer do they need to buy to ensure each employee has one?
1
Expert's answer
2013-03-29T09:00:21-0400
The abc company has 80 employees - they make 100%. So that 5% of them make 4 employees, making the total of 84 workers.

If the abc company has 1/4 computers than employees(so 4 people work on the single one) they need to buy only 4*1/4=1computer.

And if they already lack 3/4 of computers they need to buy 3/4*80+4=60+4=64 computers more.